A. During the year, L&M Leather Goods sold 1,000,000 reversible belts under a new sales promotional program. Each belt carried one rebate certificate which entitles the customer to a $4.00 cash refund when the rebate is submitted for redemption. L&M estimates that 70% of the rebates will be redeemed. 500,000 rebates had been submitted for redemption during the year. At December 31, L&M should report a liability for unredeemed rebate certificates of:

B. At the beginning of 2018, Angel Corporation began offering a two-year warranty on its products. The warranty program was expected to cost Angel 4% of net sales. Net sales made under warranty in 2018 were $180 million. Fifteen percent of the units sold were returned in 2018 and repaired or replaced at a cost of $5.3 million. The amount of warranty expense on Angel's 2018 income statement is:
Answer #1

Answer

A )

total coupons estimated to be redeemed = 1000000 * 70 % = 700000

dollar amount estimated to be redeemed = 700000 * 4 = 2800000

redeemed so far = 500000 * 4 = 2000000

800000

B )

Net Sales made under warranty in year 2018 =   $ 180 Millions

Warranty cost 4% on net sales

Warranty Expenses = Net sales * Warranty cost

                                    = $180 * 4%

                                   =   $180 * 0.04

Warranty Expenses = $7.2 Millions

So the amount of warranty Expenses on Angel's 2018 Income statement is $7.2 Millions